Booster update:
Source A and B switch, source A is not up to snuff, in LINAC, tuning beam.
Wed - 2 hr shutdown, may be longer to allow cooling.
next week longer access period, maybe April 6?

Horn update:
Ramp down BNB before noon on Wed to let it cool. 
Pressure transducer needs to be replaced next week in long shut down. (Same issue from the beginning of the experiment.)
Horn 4: aluminum seal being machined here and fitted this week.

ANNIE:
Electronics in hall, working on ORC to test phototubes. Filling tank with water this week/next week.

MicroBooNE:
Smooth running. Andy will take over next week.
Work planned for next week.